OpenStudy (anonymous):

determine the value of x^2/3 when x=8

OpenStudy (anonymous):

You can substitute 8 for x to obtain: 8^2/3 = 8*8/3 = 64/3

OpenStudy (anonymous):

are u sure u don't have to do something with the square root symbol

OpenStudy (anonymous):

Unless you actually mean x^(2/3) where the exponent is 2/3 in that case you do \[(x^{1/3})^2 = (8^{1/3})^2 = (2)^2 = 4\]

OpenStudy (anonymous):

4 is correct \[\sqrt[3]{8^2} or (\sqrt[3]{8})^2\]

OpenStudy (anonymous):

@mrjoselson yes thts how my teacher did it in class i was just confused at how @DoritosHero did it

OpenStudy (anonymous):

Yes, I'm sorry that I got confused when I thought you meant (1/3)*(x^2)

OpenStudy (anonymous):

its ok thank you

OpenStudy (anonymous):

thanks to both of u